Suppose a die is rolled twice and let
A={first toss is a prime} B={second toss is even}
Find the requested probability.(enter the probability as a fraction)
P(A n B)

Answer by math_tutor2020(3816)


A={first toss is a prime}
B={second toss is even}

set of prime numbers between 1 and 6 = {2, 3, 5}
set of even numbers between 1 and 6 = {2, 4, 6}

P(A) = probability that the first toss is a prime
P(A) = (number of primes between 1 and 6)/(number of faces on the die)
P(A) = 3/6
P(A) = 1/2

P(B) = (number of even numbers between 1 and 6)/(number of faces on the die)
P(B) = 3/6
P(B) = 1/2

P(A n B) = probability both A and B occur
P(A n B) = P(A)*P(B) which works since the events are independent
P(A n B) = (1/2)*(1/2)
P(A n B) = 1/4